# Simplex Deep Linear Discriminant Analysis

### v1

@@ Abstract
We revisit Deep Linear Discriminant Analysis (Deep LDA) from a likelihood-based perspective. While classical LDA is a simple Gaussian model with linear decision boundaries, attaching an LDA head to a neural encoder raises the question of how to train the resulting deep classifier by maximum likelihood estimation (MLE). We first show that end-to-end MLE training of an unconstrained Deep LDA model ignores discrimination: when both the LDA parameters and the encoder parameters are learned jointly, the likelihood admits a degenerate solution in which some of the class clusters may heavily overlap or even collapse, and classification performance deteriorates. Batchwise moment re-estimation of the LDA parameters does not remove this failure mode. We then propose a constrained Deep LDA formulation that fixes the class means to the vertices of a regular simplex in the latent space and restricts the shared covariance to be spherical, leaving only the priors and a single variance parameter to be learned along with the encoder. Under these geometric constraints, MLE becomes stable and yields well-separated class clusters in the latent space. On images (Fashion-MNIST, CIFAR-10, CIFAR-100), the resulting Deep LDA models achieve accuracy competitive with softmax baselines while offering a simple, interpretable latent geometry that is clearly visible in two-dimensional projections.
@@ 1 Introduction
Linear Discriminant Analysis (LDA) traces its roots to the classical works of fisher1936use and rao1948. It remains one of the most widely used classification methods and is valued for its interpretability. LDA assumes that the class label $Y\in\{1,\dots,C\}$ follows a categorical prior with probabilities $\pi=(\pi_{1},\dots,\pi_{C})$ on the probability simplex:
@@
To address this, Deep LDA [DBLP:journals/corr/DorferKW15] couples a learnable representation with an LDA head. Concretely, one first maps inputs into a latent space via a neural encoder $Z=f_{\psi}(X)$ parameterized by $\psi$. In that latent space, the pair $(Z,Y)$ is modeled according to the LDA assumptions in (1)–(2) with class means $\{\mu_{c}\}$ and a shared covariance $\Sigma$. In effect, Deep LDA replaces the conventional softmax layer with an LDA head, while letting the encoder learn a representation in which the LDA model is appropriate.
In contrast to a maximum-likelihood treatment, Deep LDA [DBLP:journals/corr/DorferKW15] optimizes Fisher’s generalized eigenvalue objective in the latent space. A straightforward maximization, however, turned out to be unstable: as noted by DBLP:journals/corr/DorferKW15, the network tends to inflate a single largest eigenvalue and essentially ignore poorly separated classes, which leads to trivial solutions. To mitigate this, they introduce a heuristic variant of the criterion that focuses only on the eigen-directions with small discriminative power (see their Eq. (9)), departing from the classical Fisher objective.
@@ 2.1 Warm-up on Classical LDA
We maximize the log-likelihood $\mathcal{L}(\theta)$ with Adam [DBLP:journals/corr/KingmaB14], treating its negative $\mathcal{L}(\theta)$ as a loss. In practice it is convenient to use unconstrained reparameterizations: (i) logits $\alpha\in\mathbb{R}^{C}$ with $\pi=\mathrm{softmax}(\alpha)$, and (ii) a Cholesky factorization $\Sigma=LL^{\top}$ with $L$ lower triangular and positive diagonal. This enforces $\pi\in\Delta^{C-1}$ and $\Sigma\succ 0$ while enabling standard auto-differentiation.
@@ 2.2 A Cautionary Experiment: Likelihood Training of Deep LDA
toward zero, while $\Sigma$ simultaneously shrinks toward near-singularity. Both effects inflate the Gaussian term $\phi(f_{\psi}(x);\mu_{y},\Sigma)$, yielding a high training likelihood. [^1] We note that replacing the log-likelihood with Fisher’s criterion does not alleviate this issue: in our experiments, optimizing the Fisher objective produced the same collapse of embeddings and class means as the maximum-likelihood formulation. A similar observation was made by DBLP:journals/corr/DorferKW15.
@@ 3 Successful Training of Deep LDA with MLE
A closely related geometric structure has been observed empirically in deep classifiers trained with cross-entropy. In particular, papyan2020prevalence document the neural collapse phenomenon: in the terminal phase of training, the last-layer features collapse to their class means, these means themselves converge to the vertices of a simplex equiangular tight frame, and the classifier weights align with those means so that prediction effectively reduces to a nearest-class-center rule. Our simplex-constrained Deep LDA head can be viewed as an explicit generative realization of a similar latent geometry: the simplex structure is imposed a priori through the class-conditional model, and the encoder is trained by maximum likelihood to populate these fixed prototypes.
@@
Training Procedure
is detailed in Algorithm 1. Unlike the batchwise re-estimation scheme in Section 2.3, there is no moment-matching: head parameters are learned purely by gradient updates.
@@ Results.
Input: Training set $\{(x_{i},y_{i})\}_{i=1}^{n}$; encoder $f_{\psi}$.
Output: Encoder parameters $\psi$; LDA-head parameters (priors $\{\pi_{c}\}$ and variance $\sigma^{2}$).
1exInitialize:
  • Construct a regular simplex in $\mathbb{R}^{C-1}$ and fix its vertices as class means $\mu_{1},\dots,\mu_{C}$ (centered at 0 and rescaled to a target pairwise distance $s$).
  • Initialize encoder parameters $\psi$ randomly.
  • Initialize priors $\pi_{c}\leftarrow 1/C$, $c=1,\ldots,C$, and variance $\sigma^{2}\leftarrow 1$.
for each epoch do
 for each minibatch $\mathcal{B}=\{(x_{j},y_{j})\}_{j=1}^{m}$ do
      Compute embeddings $z_{j}\leftarrow f_{\psi}(x_{j})$ for all $j\in\mathcal{B}$.
      Compute the minibatch negative log-likelihood
$\mathcal{J}_{\mathcal{B}}=-\frac{1}{m}\sum_{j=1}^{m}\left(\log\pi_{y_{j}}-\frac{\|z_{j}-\mu_{y_{j}}\|^{2}}{2\sigma^{2}}-\frac{d}{2}\log\sigma^{2}\right).$
Backpropagate the gradient of $\mathcal{J}_{\mathcal{B}}$ w.r.t. $\psi$, the priors $\{\pi_{c}\}$, and the variance $\sigma^{2}$. Update them with Adam (or another optimizer).
Algorithm 1 Deep LDA training with simplex means and spherical covariance (MLE)
@@
4 Experiments on Image Classification
@@
We evaluate on three standard image classification benchmarks. Fashion-MNIST [xiao2017fashion] contains 70 000 grayscale images of size $28\times 28$ from 10 classes, split into 60 000 training and 10 000 test samples. CIFAR-10 and CIFAR-100 [krizhevsky2009learning] consist of $\times32$ color images with 10 and 100 classes respectively, each with 50 000 training and 10 000 test images. Following common practice, we apply standard data augmentation to CIFAR-10/100 (random horizontal flips and random crops with padding) and use only per-pixel normalization on Fashion-MNIST.
@@ Results.
In the main experiments we set the latent dimension to d=C1d=C-1, which is the smallest dimension that can accommodate a (C1)(C-1)-dimensional simplex. To check the effect of the latent dimension, we ran an additional experiment on CIFAR-100 by varying $d\in\{150,200,\dots,500\}$. Figure 8 reports test accuracy as a function of $d$. Softmax accuracy is almost insensitive to the choice of $d$, whereas Deep LDA exhibits a mild decreasing trend as $d$ increases: enlarging the latent space offers no benefit and can slightly hurt performance. Overall, both heads remain within a narrow accuracy band for all tested dimensions, and Deep LDA is competitive with softmax for moderate $d$. This suggests that the generative head actually prefers compact latent spaces and supports our use of the minimal dimension d=C1d=C-1 as a parsimonious and empirically favorable setting.
@@ 5 Related Work
Classical Linear Discriminant Analysis dates back to fisher1936use and rao1948, who modeled each class as a Gaussian with a shared covariance matrix, yielding linear decision boundaries. Adapting LDA to deep networks has inspired several research directions, most of which modify the classical objective rather than preserve its generative form.
@@
DeepLDA [DBLP:journals/corr/DorferKW15] optimizes Fisher’s generalized eigenvalue criterion inside a neural network, with related extensions for person re-identification [DBLP:journals/pr/WuSH17] and regularized scatter objectives [chen2021regularized]. In its naive form, maximizing the mean eigenvalue of the LDA problem causes the network to favor trivial solutions that magnify a single dominant eigenvalue while neglecting poorly separated classes; this behavior was explicitly observed by Dorfer et al., who therefore replace the original objective with a heuristic that maximizes only the $k$ smallest eigenvalues (their Eqs. (8)–(9)). As a result, these methods promote between/within-class separation but no longer correspond to maximum likelihood under a generative LDA model, in contrast to our constrained Deep LDA formulation.
Discriminant losses without eigenproblems.
Other methods replace Fisher’s criterion with simpler losses that directly penalize within-class variance or enlarge between-class separation [DBLP:conf/ciss/LiDL19]. Subclass-based and center-based variants fall into this category; all reshape the latent geometry but do not model class-conditional densities.
LDA-inspired hashing and prototypes.
LDA-style structure has also appeared in hashing [DBLP:journals/mta/YanLWYCL19] and in prototypical or Gaussian-based classifiers. These prototype methods often use distances as logits trained with cross-entropy, making them discriminative rather than generative and lacking explicit likelihood-based training.
@@
Prior deep LDA variants either modify Fisher’s criterion or use Gaussian distances in a discriminative loss. In contrast, we retain the generative LDA model and train it by maximum likelihood, made stable by constraining the class means to a fixed simplex and enforcing a spherical shared covariance. These constraints remove the degeneracies of unconstrained deep LDA (Section 2) and yield an interpretable, well-behaved generative classifier.
@@ 6 Conclusion
We revisited Deep Linear Discriminant Analysis from a likelihood perspective and showed that the unconstrained deep model is ill-posed under maximum-likelihood training: the encoder collapses embeddings, some of the class means drift toward each other, and the shared covariance shrinks toward singularity, yielding high likelihood but poor discrimination. A simple geometric constraint—fixing the class means to a regular simplex and enforcing a spherical shared covariance—removes the pathological tendency of the class clusters to collapse or heavily overlap, and improves the classification performance. On synthetic data and image benchmarks, the resulting Deep LDA classifier yields clean, interpretable latent spaces with well-separated clusters and accuracy competitive with softmax, illustrating that mild geometric structure can restore the viability of likelihood-based training for deep discriminant analysis.

### v2

@@ Abstract
We revisit Deep Linear Discriminant Analysis (Deep LDA) from a likelihood-based perspective. While classical LDA is a simple Gaussian model with linear decision boundaries, attaching an LDA head to a neural encoder raises the question of how to train the resulting deep classifier by maximum likelihood estimation (MLE). We first show that end-to-end MLE training of an unconstrained Deep LDA model ignores discrimination: when both the LDA parameters and the encoder parameters are learned jointly, the likelihood admits a degenerate solution in which some of the class clusters may heavily overlap or even collapse, and classification performance deteriorates. Batchwise moment re-estimation of the LDA parameters does not remove this failure mode. We then propose a constrained Deep LDA formulation that fixes the class means to the vertices of a regular simplex in the latent space and restricts the shared covariance to be spherical, leaving only the priors and a single variance parameter to be learned along with the encoder. Under these geometric constraints, MLE becomes stable and yields well-separated class clusters in the latent space. On images (Fashion-MNIST, CIFAR-10, CIFAR-100) and texts (AG News, CLINC150), the resulting Deep LDA models achieve accuracy competitive with softmax baselines while offering a simple, interpretable latent geometry that is clearly visible in two-dimensional projections.
@@ 1 Introduction
Linear Discriminant Analysis (LDA) traces its roots to the classical works of Fisher [4] and Rao [12]. It remains one of the most widely used classification methods and is valued for its interpretability. LDA assumes that the class label $Y\in\{1,\dots,C\}$ follows a categorical prior with probabilities $\pi=(\pi_{1},\dots,\pi_{C})$ on the probability simplex:
@@
To address this, Deep LDA [3] couples a learnable representation with an LDA head. Concretely, one first maps inputs into a latent space via a neural encoder $Z=f_{\psi}(X)$ parameterized by $\psi$. In that latent space, the pair $(Z,Y)$ is modeled according to the LDA assumptions in (1)–(2) with class means $\{\mu_{c}\}$ and a shared covariance $\Sigma$. In effect, Deep LDA replaces the conventional softmax layer with an LDA head, while letting the encoder learn a representation in which the LDA model is appropriate.
In contrast to a maximum-likelihood treatment, Deep LDA [3] optimizes Fisher’s generalized eigenvalue objective in the latent space. A straightforward maximization, however, turned out to be unstable: as noted by Dorfer et al. [3], the network tends to inflate a single largest eigenvalue and essentially ignore poorly separated classes, which leads to trivial solutions. To mitigate this, they introduce a heuristic variant of the criterion that focuses only on the eigen-directions with small discriminative power (see their Eq. (9)), departing from the classical Fisher objective.
@@ 2.1 Warm-up on Classical LDA
We maximize the log-likelihood $\mathcal{L}(\theta)$ with Adam, treating its negative $\mathcal{L}(\theta)$ as a loss. In practice it is convenient to use unconstrained reparameterizations: (i) logits $\alpha\in\mathbb{R}^{C}$ with $\pi=\mathrm{softmax}(\alpha)$, and (ii) a Cholesky factorization $\Sigma=LL^{\top}$ with $L$ lower triangular and positive diagonal. This enforces $\pi\in\Delta^{C-1}$ and $\Sigma\succ 0$ while enabling standard auto-differentiation.
@@ 2.2 A Cautionary Experiment: Likelihood Training of Deep LDA
toward zero, while $\Sigma$ simultaneously shrinks toward near-singularity. Both effects inflate the Gaussian term $\phi(f_{\psi}(x);\mu_{y},\Sigma)$, yielding a high training likelihood. [^1] We note that replacing the log-likelihood with Fisher’s criterion does not alleviate this issue: in our experiments, optimizing the Fisher objective produced the same collapse of embeddings and class means as the maximum-likelihood formulation. A similar observation was made by Dorfer et al. [3].
@@ 3 Successful Training of Deep LDA with MLE
A closely related geometric structure has been observed empirically in deep classifiers trained with cross-entropy. In particular, Papyan et al. [10] document the neural collapse phenomenon: in the terminal phase of training, the last-layer features collapse to their class means, these means themselves converge to the vertices of a simplex equiangular tight frame, and the classifier weights align with those means so that prediction effectively reduces to a nearest-class-center rule. Our simplex-constrained Deep LDA head can be viewed as an explicit generative realization of a similar latent geometry: the simplex structure is imposed a priori through the class-conditional model, and the encoder is trained by maximum likelihood to populate these fixed prototypes.
@@
Why maximum likelihood.
A central goal of this work is to train the Deep LDA head by maximum likelihood (equivalently, minimizing negative log-likelihood). Prior “Deep LDA” style approaches typically optimize modified Fisher-type criteria rather than the LDA likelihood directly [3, 13, 1], or use discriminative objectives defined over Gaussian/Mahalanobis distances [9]. We make this distinction explicit because our contribution is precisely to obtain a well-posed likelihood-based formulation under simple geometric constraints.
Training
follows standard end-to-end maximum-likelihood optimization: we fix the simplex means $\{\mu_{c}\}$, and jointly update the encoder parameters $\psi$ and head parameters $(\pi,\sigma^{2})$ by minimizing the negative log-likelihood (7) with Adam. Unlike the batchwise re-estimation scheme in Section 2.3, there is no moment-matching: all learnable parameters are updated by backpropagation. Our implementation is available at https://github.com/zh3nis/simplex-lda.
@@ Results.
4 Experiments on Real Data
@@
We evaluate on three standard image classification benchmarks. Fashion-MNIST [14] contains 70 000 grayscale images of size $28\times 28$ from 10 classes, split into 60 000 training and 10 000 test samples. CIFAR-10 and CIFAR-100 [6] consist of $\times32$ color images with 10 and 100 classes respectively, each with 50 000 training and 10 000 test images. Following common practice, we apply standard data augmentation to CIFAR-10/100 (random horizontal flips and random crops with padding) and use only per-pixel normalization on Fashion-MNIST.
@@ Results.
In the main experiments we set the latent dimension to d=C1d=C-1, which is the smallest dimension that can accommodate a (C1)(C-1)-dimensional simplex. To check the effect of the latent dimension, we ran an additional experiment on CIFAR-100 by varying $d\in\{150,200,\dots,500\}$. Figure 8 reports test accuracy as a function of $d$. Softmax accuracy is almost insensitive to the choice of $d$, whereas Deep LDA exhibits a mild decreasing trend as $d$ increases: enlarging the latent space offers no benefit and can slightly hurt performance.
@@
Overall, both heads remain within a narrow accuracy band for all tested dimensions, and Deep LDA is competitive with softmax for moderate $d$. This suggests that the generative head actually prefers compact latent spaces and supports our use of the minimal dimension d=C1d=C-1 as a parsimonious and empirically favorable setting.
Covariance structure ablation.
Our default head uses a spherical shared covariance, $\Sigma=\sigma^{2}I$, to keep the head lightweight and comparable in parameter count to a linear softmax layer at the same embedding dimension. To test whether additional covariance flexibility helps, we train simplex-based Deep LDA with spherical, diagonal, and full covariances on FashionMNIST, CIFAR-10, and CIFAR-100 (3 runs per setting). Table 2 reports mean $\pm$ std test accuracy. Neither diagonal nor full covariance improves performance; in fact, full covariance significantly degrades accuracy on CIFAR-100. This suggests that the encoder can adapt to a spherical covariance structure, while additional covariance degrees of freedom may introduce optimization/estimation difficulties.
Table 2: Ablation of the shared covariance parameterization in simplex-based Deep LDA (mean $\pm$ std over 3 runs).
Covariance FashionMNIST CIFAR-10 CIFAR-100
Spherical $93.38\pm 0.06$ $90.49\pm 0.26$ $65.20\pm 0.22$
Diagonal $93.32\pm 0.07$ $90.31\pm 0.26$ $65.02\pm 0.09$
Full $93.31\pm 0.11$ $89.91\pm 0.31$ $59.64\pm 0.71$
Stronger off-the-shelf backbone (ResNet-50).
We additionally evaluate our approach with a stronger, widely used backbone. Specifically, we fine-tune an ImageNet-pretrained ResNet-50 encoder and compare (i) our Simplex LDA likelihood-trained head against (ii) a standard Softmax head trained with cross-entropy. To ensure a controlled comparison, we use the same encoder, the same embedding dimension for both heads, setting D=C1D=C-1. Table 3 reports results over 3 runs (mean $\pm$ std). Overall, with a substantially stronger backbone, the simplex-constrained likelihood-trained head remains competitive with the conventional softmax baseline across datasets.
Table 3: Fine-tuning a pretrained ResNet-50 backbone (mean $\pm$ std over 3 runs).
Head FashionMNIST CIFAR-10 CIFAR-100
Softmax $93.99\pm 0.17$ $96.30\pm 0.14$ $80.85\pm 0.19$
Simplex LDA $94.22\pm 0.08$ $96.35\pm 0.19$ $80.45\pm 0.22$
Beyond images: text classification with a pretrained Transformer.
To demonstrate that the proposed head generalizes beyond image benchmarks, we fine-tune a pretrained Transformer (BERT-base-uncased [2]) for text classification on AG News (4 classes; 120k train / 7.6k test) [^4] http://groups.di.unipi.it/~gulli/AG_corpus_of_news_articles.html and CLINC150 (151 classes; 15k train / 5.5k test) [7]. We compare (i) a standard linear Softmax head to (ii) our SimplexLDA head, using an architecture-matched setup with an embedding projection to d=C1d=C-1. Both models are fine-tuned with AdamW (learning rate $2\times 10^{-5}$, batch size 32) for 3 epochs, and results are averaged over 3 random seeds. Table 4 reports test accuracy (mean $\pm$ std). Overall, SimplexLDA remains competitive with Softmax on AG News and improves performance on CLINC150, supporting the generality and scalability of the proposed likelihood-trained head.
Table 4: Text classification: test accuracy (mean $\pm$ std over 3 runs).
Head AG News CLINC150
Softmax $94.40\pm 0.07$ $87.11\pm 0.56$
SimplexLDA $94.37\pm 0.37$ $88.12\pm 0.34$
@@ 5 Related Work
Classical Linear Discriminant Analysis dates back to Fisher [4] and Rao [12], who modeled each class as a Gaussian with a shared covariance matrix, yielding linear decision boundaries. Adapting LDA to deep networks has inspired several research directions, most of which modify the classical objective rather than preserve its generative form.
@@
DeepLDA [3] optimizes Fisher-style generalized eigenvalue criterion, with related extensions for person re-identification [13] and regularized scatter objectives [1]. In its naive form, maximizing the mean eigenvalue of the LDA problem causes the network to favor trivial solutions that magnify a single dominant eigenvalue while neglecting poorly separated classes; this behavior was explicitly observed by Dorfer et al. [3], who therefore replace the original objective with a heuristic that maximizes only the $k$ smallest eigenvalues (their Eqs. (8)–(9)). As a result, these methods promote between/within-class separation but no longer correspond to maximum likelihood under a generative LDA model, in contrast to our constrained Deep LDA formulation.
Fixed classifiers and simplex-structured heads.
A complementary line of work studies fixing the final classifier layer to impose a prescribed geometry. Hoffer et al. [5] argue that the last linear layer can be fixed (up to scale) with little loss in accuracy. Building on this idea, Pernici et al. [11] fix class vectors to the vertices of regular polytopes (including the simplex) to obtain maximally separated and stationary embeddings. Motivated by the neural collapse phenomenon [10], several works explicitly use a regular simplex (simplex equiangular tight frame) structure for the classifier and/or class centers, either by analyzing the induced optimization landscape [16], by demonstrating that fixing a simplex ETF classifier can induce neural collapse even under class imbalance [15], or by proposing to optimize towards the nearest simplex ETF rather than hard-fixing one for improved stability and convergence [8]. Importantly, in these works the simplex (or polytope) structure is primarily leveraged to improve discriminative training with losses such as cross-entropy or squared error.
@@
Prior deep LDA variants either modify Fisher’s criterion or use Gaussian distances in a discriminative loss, and fixed/simplex classifier works use simplex structure mainly as an inductive bias for discriminative training. In contrast, we retain the generative LDA model and train it by maximum likelihood. In our setting, the regular simplex plays a different role: it acts as a natural constraint that makes end-to-end likelihood training of Deep LDA well-posed, preventing the degeneracies that arise under an unconstrained model (Section 2). Together with a spherical shared covariance, these constraints yield an interpretable, well-behaved generative classifier whose MLE training produces separated class clusters without collapse.
@@ 6 Conclusion
We revisited Deep Linear Discriminant Analysis from a likelihood perspective and showed that the unconstrained deep model is ill-posed under maximum-likelihood training: the encoder collapses embeddings, some of the class means drift toward each other, and the shared covariance shrinks toward singularity, yielding high likelihood but poor discrimination. A simple geometric constraint—fixing the class means to a regular simplex and enforcing a spherical shared covariance—removes the pathological tendency of the class clusters to collapse or heavily overlap, and improves the classification performance. On synthetic data and image benchmarks, the resulting Deep LDA classifier yields clean, interpretable latent spaces with well-separated clusters and accuracy competitive with softmax, illustrating that mild geometric structure can restore the viability of likelihood-based training for deep discriminant analysis. We additionally verified that the head transfers beyond images by fine-tuning a pretrained Transformer on text classification (AG News, CLINC150), where it remains competitive with Softmax.
Acknowledgements
This work was supported by the Science Committee of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP26198325).